Angles E and F are complimentary.
If the measure of angle E = x-10 and the measure of angle F = x+2, find the value of x and the measure of each angle.

Answer by LinnW(1048)

Since E and F are complementary
E + F = 90
(x-10) + (x+2) = 90
2x -8 = 90
add 8 to each side
2x = 98
divide each side by 2
x = 49
E = x-10 = 49 - 10
E = 39
F = x+2 = 49 + 2
F = 51